# Towards Precise 3D Quantum Control of a Levitated Dipolar Scatterer using Spatial Mode Decomposition

### v1

@@ III Deriving the Measurement Imprecision
where we have discarded the term proportional to $q(t)^{2}$ under the argument that it contributes only weakly to the measurement result. We can use the form of the recovered power to estimate the measurement imprecision, $S_{\mathrm{imp}}$ [15]. This involves first collecting those terms independent of the nanoparticle’s motion, which we identify as contributing only noise.
@@
$S^{\ell,m}_{\mathrm{noise}}$ $=\frac{\hbar kc}{2\pi}\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\Big{[}(\eta^{\ell,m}_{\mathrm{ref}})^{2}\iint_{\mathrm{FMF}}|\bm{E}_{\mathrm{ref}}(x,y)|^{2}~{}\mathrm{d}x\mathrm{d}y+(\eta^{\ell,m}_{0})^{2}\iint_{\mathrm{FMF}}|\bm{E}_{0}(x,y)|^{2}~{}\mathrm{d}x\mathrm{d}y$
$\hskip 128.0374pt+2\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{0}\iint_{\mathrm{FMF}}|\bm{E}_{\mathrm{ref}}^{*}(x,y)\bm{E}_{0}(x,y)|~{}\mathrm{d}x\mathrm{d}y\Big{]}~{}.$ (S19)
Additionally, in order to extract $q(t)$, the recovered power must be weighted by a factor
$\beta_{q}^{\ell,m}$ $=\frac{2}{w_{0}}\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\Big{[}\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{q}\iint_{\mathrm{FMF}}|\bm{E}_{\mathrm{ref}}^{*}(x,y)\bm{E}^{(q)}(x,y)|~{}\mathrm{d}x\mathrm{d}y$
$\hskip 128.0374pt+\eta^{\ell,m}_{0}\eta^{\ell,m}_{q}\iint_{\mathrm{FMF}}|\bm{E}_{0}^{*}(x,y)\bm{E}^{(q)}(x,y)|~{}\mathrm{d}x\mathrm{d}y\Big{]}~{},$ (S20)
which is subsequently squared in taking the power spectral density. The measurement imprecision associated with the nanoparticle’s motion along $\bm{\hat{e}}_{q}$, in the LPℓ,m channel, will therefore be of the form $S_{\mathrm{imp},q}^{\ell,m}=S^{\ell,m}_{\mathrm{noise}}~{}/~{}(\beta_{q}^{\ell,m})^{2}$.
Let us now consider how one might go about minimising this value. Recognising that the reference field is often the dominant source of shot noise, given $|\bm{E}_{\mathrm{ref}}(x,y)|~{}\gg~{}|\bm{E}_{\mathrm{scat}}(x,y~{};\bm{q}(t))|$, we opt to consider the case wherein it is effectively de-coupled from each LP mode ($\eta^{\ell,m}_{\mathrm{ref}}=0$). This would be considered disadvantageous in conventional detection systems, where the reference field interferometrically extracts information stored in the phase of the scattered light. Here, however, information is instead encoded in amplitude modulations of the LP modes upon coupling into the photonic chip. In this case, the measurement imprecision reduces to
$S^{\ell,m}_{\mathrm{imp},q}$ $=\frac{\hbar kc}{8\pi}\sqrt{\frac{\mu_{0}}{\varepsilon_{0}}}\frac{\iint_{\mathrm{FMF}}|\bm{E}_{0}(x,y)|^{2}~{}\mathrm{d}x\mathrm{d}y}{\left(\iint_{\mathrm{FMF}}|\bm{E}_{0}^{*}(x,y)\bm{E}^{(q)}(x,y)|~{}\mathrm{d}x\mathrm{d}y\right)^{2}}\frac{w_{0}^{2}}{(\eta^{\ell,m}_{q})^{2}}~{},$ (S21)
@@
which is minimised when $\eta^{\ell,m}_{q}=1$. Perhaps counter-intuitively, this result suggests that the measurement imprecision is independent of our coupling to the elastically scattered field. An identical result is found when we consider de-coupling from the elastically scattered field, where the resultant measurement imprecision is independent of our coupling to the reference field.
IV Discussion of Losses
@@
Here, we provide a brief overview of the different efficiencies reported in the main text.
IV.1 Collection Efficiency
We write the effective numerical aperture (NA) of the parabolic mirror used here as
$\mathrm{NA}$ $=1-\mathrm{cos}\left(\mathrm{arctan}\left[\frac{r_{\mathrm{m}}}{f_{\mathrm{m}}-\frac{r_{\mathrm{m}}^{2}}{4f_{\mathrm{m}}}}\right]\right)~{},$ (S22)
@@ IV.1 Collection Efficiency
where $f_{\mathrm{m}}$ and $r_{\mathrm{m}}$ are the focal length and radius of the parabolic mirror, respectively. For the design specifications used, this implies $\mathrm{NA}=1$. Hence, we expect to collect all of the light back-scattered from an optically-levitated nanoparticle, equivalent to a photon collection efficiency of $\eta^{*}=0.5$ for all three center-of-mass (COM) translational modes. In the context of information losses, however, we require a more nuanced approach. This is because motional information is not uniformly distributed in the scattered field. To this end, we model the so-called information radiation patterns associated with each COM translational mode, $q(t)$ for $q=x,y,z$ [15, 14]. These are of the form
$\mathcal{I}^{(x)}(\theta,\varphi)$ $=\frac{15}{8\pi}\Big{[}1-\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\phi)\Big{]}~{}\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\varphi)~{}\mathrm{d}\Omega$ (S23)
$\mathcal{I}^{(y)}(\theta,\varphi)$ $=\frac{15}{16\pi}\Big{[}1-\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\phi)\Big{]}~{}\mathrm{sin}^{2}(\theta)\mathrm{sin}^{2}(\varphi)~{}\mathrm{d}\Omega$ (S24)
$\mathcal{I}^{(z)}(\theta,\varphi)$ $=\frac{3}{8\pi(\frac{2}{5}+A^{2})}\Big{[}1-\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\phi)\Big{]}\Big{(}\mathrm{cos}(\theta)-A\Big{)}^{2}~{}\mathrm{d}\Omega~{},$ (S25)
@@
for an $x$-polarised dipolar scatterer. Here, $A$ is a consequence of the Gouy phase inherited from the trapping beam. Note that, for intermediate values of NA, we find $A\approx 1-(kz_{\mathrm{R}})^{-1}$ for Rayleigh length $z_{\mathrm{R}}$. Given our setup operates well beyond this limit, however, we choose to focus on illumination by a plane wave ($A=1$). To determine the information collection efficiency for the parabolic mirror, we evaluate the integral
$\eta_{q}$ $=\int^{2\pi}_{0}\int^{\pi}_{\pi/2}\mathcal{I}^{(q)}(\theta,\varphi)~{}\mathrm{d}\theta\mathrm{d}\varphi~{},$ (S26)
@@
where the lower integration limit on $\theta$ is given by $\pi-\mathrm{arcsin}(\mathrm{NA})$. This represents the amount of information content collected in the back focal plane of the optical trap, and gives $(\eta_{x},\eta_{y},\eta_{z})=(0.50,0.50,0.90)$, respectively.
IV.2 Path Losses
@@
To estimate the path losses, we measure the difference in power of the forward- and backward-propagating trapping beam in the focal plane of the collection lens, with no nanoparticle trapped. From this, we arrive at a value for the path losses of $0.10$, implying an efficiency of $0.90$.
IV.3 Coupling Efficiency
@@
IV.4 Quantum Efficiency
From the technical specifications provided by the supplier, we estimate the quantum efficiency of our detectors to be $0.72$.
IV.5 Dark Noise
Additional information loss must be considered in the electronic line, after each signal has been measured. In particular, we note that the so-called dark current noise will mask some of the available information. At frequencies near to the natural oscillation frequencies of the nanoparticle, we measure the dark noise to be approximately $8$ dB below the shot noise, corresponding to an efficiency of $0.79$.
IV.6 Digital Noise
Finally, in the context of measurement-based control techniques, it is meaningful to consider those additional losses which stem from use of a digital feedback loop. In particular, here we are interested in using a Field Programmable Gate Array (FPGA) (Red Pitaya board) which functions as a phase-locked loop (PLL), producing a modulated signal to be used in Parametric Feedback Cooling (PFC). This circuit induces an effective information loss of $0.95$.
V Feedback Cooling Data
Figure S1: Calibrated power spectral density (PSD) $S_{qq}$ of the nanoparticle’s COM motion at thermal equilibrium (300 $\mathrm{K}$) and after having been cooled to a temperature of $(T_{x},T_{y},T_{z})=(100,7,20)$ mK at $4.7\times 10^{-5}$ mbar via parametric feedback cooling.
VI Calculation of Calibration Factor
In the limit of small mechanical displacements, the thermally-driven center-of-mass (COM) motion of an optically-levitated dielectric nanoparticle can be approximated as simple harmonic motion. Specifically, we are interested in the limit where each of the three spatial dimensions are decoupled from one another, and so may be treated separately. Hence, in the presence of a feedback force, we write
$\ddot{q}(t)+\left[\gamma+\delta\gamma^{(q)}\right]\dot{q}(t)+\left[\Omega_{q}+\delta\Omega_{q}\right]^{2}q(t)$ $=\frac{1}{m}\bm{F}_{\mathrm{th}}^{(q)}(t)~{},$ (S27)
@@ VI Calculation of Calibration Factor
$\langle\bm{F}_{\mathrm{th}}(t)\bm{F}_{\mathrm{th}}(t+~{}\tau)\rangle=2m\gamma k_{B}T_{0}~{}\delta(t-\tau)~{}.$ (S28)
@@
Note that, in equation (S27) the feedback force is encoded in the linear shifts $\delta\gamma^{(q)}$ and $\delta\Omega_{q}$, respectively. From this equation of motion, we can construct a power spectral density (PSD) for this system of the form
$S_{(qq)}(\Omega)$ $=\chi^{2}\frac{k_{B}T_{0}}{\pi m}\frac{\gamma}{\left([\Omega_{q}+\delta\Omega_{q}]^{2}-\Omega^{2}\right)^{2}+\Omega^{2}\left(\gamma+\delta\gamma^{(q)}\right)^{2}}~{},$ (S29)
@@
where $\chi$ is referred to as the calibration factor, in units of V / m. This term stems from the fact that most experimental PSDs, calculated from the time-trace recorded on a photodetector, are inherently in units of V2 / Hz. However, given we are concerned with the motion of a nanoparticle, we require a PSD in units of m2 / Hz to determine relevant quantities, such as the position resolution. For ease of fitting, here we choose to simplify this PSD to one of the form
$S_{(qq)}(\Omega)$ $=\frac{A}{(B^{2}-\Omega^{2})^{2}+\Omega^{2}C^{2}}~{}.$ (S30)
@@
Hence, by fitting equation (S30) to data taken at thermal equilibrium ($T_{0}=300$ K), we can extract the calibration factor in terms of known quantities. That is,
$\chi=\sqrt{\frac{A}{C}\frac{\pi m}{k_{B}T_{0}}}~{}.$ (S31)
VII Temperature Estimation

### v2

@@ Appendix C Deriving the measurement imprecision
where we have discarded the term proportional to $[q(t)~{}/~{}w_{0}]^{2}$ as we expect the displacements to be far smaller than $w_{0}$ along each axis. Note that, from Eq. (31), we do not simply measure $q(t)$. Instead, the measurement result is dominated by a series of time-independent coefficients, which mask the information available on the nanoparticle’s position. From the form of these coefficients, we can estimate the measurement imprecision. This involves first collecting those terms independent of the nanoparticle’s motion, which we identify as contributing only shot noise, and so have a power spectral density (PSD) of the form
@@
$S^{\ell,m}_{\mathrm{noise}}$ $=\frac{\hbar kc}{2\pi}\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\Bigg{(}(\eta^{\ell,m}_{\mathrm{ref}})^{2}\alpha_{\mathrm{ref}}+(\eta^{\ell,m}_{0})^{2}\alpha_{0}^{2}+2\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{0}\alpha_{\mathrm{ref}}\alpha_{0}\Bigg{)}~{}.$ (32)
Additionally, $q(t)$ is weighted by a factor of
$\beta_{q}^{\ell,m}$ $=\frac{2}{w_{0}}\sqrt{\frac{\varepsilon_{0}}{\mu_{0}}}\Bigg{(}\eta^{\ell,m}_{\mathrm{ref}}\eta^{\ell,m}_{q}\alpha_{\mathrm{ref}}\alpha_{q}+\eta^{\ell,m}_{0}\eta^{\ell,m}_{q}\alpha_{0}\alpha_{q}\Bigg{)}~{},$ (33)
which is squared when calculating the PSD. The imprecision associated with measuring the nanoparticle’s motion along the $\bm{\hat{e}}_{q}$ axis, in the LPℓ,m output channel of the spatial mode sorter, will therefore be of the form ${S_{\mathrm{imp},q}^{\ell,m}=S^{\ell,m}_{\mathrm{noise}}~{}/~{}(\beta_{q}^{\ell,m})^{2}}$ [14, 15]. Let us now consider how one might go about minimising this measurement imprecision. In particular, given the coupling coefficients calculated in Fig. 3 of the main text, we opt to consider the case wherein the LO field is effectively decoupled from each LP mode ($\eta^{\ell,m}_{\mathrm{ref}}=0$). In this case, the measurement imprecision reduces to
$S^{\ell,m}_{\mathrm{imp},q}$ $=\frac{\hbar kc}{8\pi}\sqrt{\frac{\mu_{0}}{\varepsilon_{0}}}\frac{1}{\alpha_{q}^{2}}\frac{w_{0}^{2}}{(\eta^{\ell,m}_{q})^{2}}~{},$ (34)
@@
which is minimised when $\eta^{\ell,m}_{q}=1$. Perhaps counter-intuitively, this result suggests that the measurement imprecision is independent of the overlap of the inelastically and elastically scattered field, despite Eq. (31) suggesting that our measurement relies on the interference of these two fields. Note that, in the case where each LP mode is instead decoupled from the elastically scattered field ($\eta^{\ell,m}_{0}=0$), we arrive at a very similar result. That is,
$S^{\ell,m}_{\mathrm{imp},q}$ $=\frac{\hbar kc}{8\pi}\sqrt{\frac{\mu_{0}}{\varepsilon_{0}}}\frac{1}{\alpha_{q}^{2}}\frac{w_{0}^{2}}{(\eta^{\ell,m}_{q})^{2}}~{},$ (35)
which is independent of the overlap of the inelastically scattered and LO fields.
Appendix D Form of the power spectral density
Let us now consider the form of the power spectral density (PSD) measured at each LPℓ,m output channel of the space-division de-multiplexing photonic chip. In Appendix C we argued that the power measured at a detector is given by Eq. (31), and can be split into those terms independent of the nanoparticle’s motion, and those terms proportional to $q(t)$. We have already identified the former as sources of noise in the measurement result, with a PSD given by Eq. (32). Collecting the remaining terms, we introduce $s_{qq}^{\ell,m}(t)=q(t)\beta_{q}^{\ell,m}$, where $\beta_{q}^{\ell,m}$ is defined in Eq. (33). From the Wiener-Khinchin theorem, this quantity has a PSD of the form
$S_{qq}^{\ell,m}(\Omega)$ $=\int^{\infty}_{-\infty}\langle s_{qq}^{\ell,m}(t)s_{qq}^{\ell,m}(t+\tau)\rangle e^{-i\Omega\tau}~{}\mathrm{d}\tau$
$=(\beta_{q}^{\ell,m})^{2}~{}\int^{\infty}_{-\infty}\langle q(t)q(t+\tau)\rangle e^{-i\Omega\tau}~{}\mathrm{d}\tau~{}.$ (36)
From the equation of motion for an optically-levitated nanoparticle, discussed in Appendix (G), the second term in this expression reduces to the well-known result
$S_{qq}^{\ell,m}(\Omega)$ $=(\beta_{q}^{\ell,m})^{2}\frac{k_{B}T_{0}}{\pi m}\frac{\gamma}{\left(\Omega_{q}^{2}-\Omega^{2}\right)^{2}+\Omega^{2}\gamma^{2}}~{}.$ (37)
The total PSD measured at each LP output channel of the space-division de-multiplexing photonic chip is then of the form
$S^{\ell,m}_{\mathrm{tot}}(\Omega)$ $=S^{\ell,m}_{\mathrm{imp},q}+S^{\ell,m}_{qq}(\Omega)~{}.$ (38)
Here we see the role of the imprecision in this measurement, derived in Appendix (C), which functions as the noise floor of the PSD.
Appendix E Discussion of losses and efficiencies
@@
Here, we provide a brief overview of the efficiencies reported in the main text.
E.1 Collection efficiency
We write the effective numerical aperture (NA) of the parabolic mirror as
$\mathrm{NA}$ $=1-\mathrm{cos}\left(\mathrm{arctan}\left[\frac{r_{\mathrm{m}}}{f_{\mathrm{m}}-\frac{r_{\mathrm{m}}^{2}}{4f_{\mathrm{m}}}}\right]\right)~{},$ (39)
@@ E.1 Collection efficiency
where $f_{\mathrm{m}}$ and $r_{\mathrm{m}}$ are the focal length (1 mm) and radius (2 mm) of the parabolic mirror, respectively. For the design specifications used here, this implies a numerical aperture of $\mathrm{NA}=1$. Hence, we expect to collect all of the light back-scattered from an optically-levitated nanoparticle, equivalent to a photon collection efficiency of $50\%$ for all three center-of-mass (COM) translational degrees of freedom (DOFs). In the context of information losses, however, we require a more nuanced approach. This is because motional information is not uniformly distributed in the scattered field. To this end, we consider the so-called information radiation patterns (IRPs) associated with each COM translational DOF, $q(t)$ for $q=x,y,z$ [14, 15]. These IRPs are of the form
$\mathcal{I}_{x}(\theta,\varphi)$ $=\frac{15}{8\pi}\Big{[}1-\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\phi)\Big{]}~{}\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\varphi)~{}\mathrm{d}\Omega$ (40)
$\mathcal{I}_{y}(\theta,\varphi)$ $=\frac{15}{16\pi}\Big{[}1-\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\phi)\Big{]}~{}\mathrm{sin}^{2}(\theta)\mathrm{sin}^{2}(\varphi)~{}\mathrm{d}\Omega$ (41)
$\mathcal{I}_{z}(\theta,\varphi)$ $=\frac{3}{8\pi(\frac{2}{5}+A^{2})}\Big{[}1-\mathrm{sin}^{2}(\theta)\mathrm{cos}^{2}(\phi)\Big{]}\Big{(}\mathrm{cos}(\theta)-A\Big{)}^{2}~{}\mathrm{d}\Omega~{},$ (42)
@@
for an $x$-polarised dipolar scatterer. Here, $A$ is a consequence of the Gouy phase inherited from the trapping beam. Note that, for intermediate values of the numerical aperture, we find $A\approx 1-(kz_{\mathrm{R}})^{-1}$ for Rayleigh length $z_{\mathrm{R}}$. Given that our setup operates well beyond this limit, however, we choose to focus on illumination by a plane wave ($A=1$). To determine the information collection efficiency for the parabolic mirror, we evaluate the integral
$\eta_{q}$ $=\int^{2\pi}_{0}\int^{\pi}_{\pi/2}\mathcal{I}_{q}(\theta,\varphi)~{}\mathrm{d}\theta\mathrm{d}\varphi~{},$ (43)
@@
where the lower integration limit on $\theta$ is given by $\pi-\mathrm{arcsin}(\mathrm{NA})$. This represents the amount of information content collected in the back focal plane of the optical trap, and gives $(\eta_{x},\eta_{y},\eta_{z})=(50,50,90)\%$, respectively.
E.2 Path losses
@@
To estimate the path losses, we measure the difference in power of the forward- and backward-propagating trapping beam in the focal plane of the collection lens, with no nanoparticle trapped. From this, we arrive at a value for the path losses of approximately $10\%$, equivalent to an efficiency of $90\%$.
E.3 Coupling efficiency
@@
E.4 Quantum efficiency
From the technical specifications provided by the supplier of the detectors used here (PD100-DC, Koheron), we estimate the quantum efficiency of our detectors to be $0.72\%$ at $\lambda=1550$ nm.
E.5 Dark noise
Information loss must also be considered in the electronic line, after each signal has been measured. In particular, we note that the dark current noise will mask some of the available information. At frequencies near to the natural oscillation frequencies of the nanoparticle, we measure the dark noise to be approximately $8$ dB below the shot noise, equivalent to an efficiency of $79\%$.
E.6 Digital noise
In the context of measurement-based control techniques, it is meaningful to consider those additional losses which stem from use of a digital feedback loop. In particular, here we are interested in using a Field Programmable Gate Array (FPGA) board (STEMlab 125-14, Red Pitaya) which functions as a phase-locked loop (PLL), with an estimated efficiency of $95\%$.
Appendix F Feedback cooling data
Figure 5: Calibrated power spectral density (PSD) $S_{qq}$ of the nanoparticle’s COM motion at thermal equilibrium (300 $\mathrm{K}$) and after having been cooled to a temperature of $(T_{x},T_{y},T_{z})=(7,106,21.4)\pm(3.1,12,2.3)$ mK at $4.7\times 10^{-5}$ mbar via parametric feedback cooling.
Here we show a comparison of the power spectral densities (PSDs) for the nanoparticle’s center-of-mass translational degrees of freedom at thermal equilibrium (300 K) (”Hot”) and after having been cooled to a temperature of ${(T_{x},T_{y},T_{z})=(7,106,21.4)\pm(3.1,12,2.3)}$ mK (”Cold”). These results were measured at a pressure of $4.7\times 10^{-5}$ mbar, which is the current technical limitation of our vacuum chamber. To calibrate each PSD and estimate the corresponding temperatures we used those methods described in Appendices G and H, respectively.
Appendix G Calculation of calibration factor
In the limit of small mechanical displacements, the thermally driven center-of-mass (COM) motion of an optically levitated dielectric nanoparticle can be approximated as simple harmonic motion. Specifically, we are interested in the limit where each of the three spatial dimensions are decoupled from one another, and so may be treated separately. Hence, in the presence of a feedback force, we write
$\ddot{q}(t)+\left[\gamma+\delta\gamma_{q}\right]\dot{q}(t)+\left[\Omega_{q}+\delta\Omega_{q}\right]^{2}q(t)$ $=\frac{1}{m}\bm{F}_{\mathrm{th}}^{(q)}(t)~{},$ (44)
@@ Appendix G Calculation of calibration factor
$\langle\bm{F}_{\mathrm{th}}(t)\bm{F}_{\mathrm{th}}(t+~{}\tau)\rangle=2m\gamma k_{B}T_{0}~{}\delta(t-\tau)~{}.$ (45)
@@
Note that, in Eq. (44) the feedback force is encoded in the linear shifts $\delta\gamma_{q}$ and $\delta\Omega_{q}$, respectively. From this equation of motion, we can construct a power spectral density (PSD) for this system of the form
$S_{(qq)}(\Omega)$ $=\chi^{2}\frac{k_{B}T_{0}}{\pi m}\frac{\gamma}{\left([\Omega_{q}+\delta\Omega_{q}]^{2}-\Omega^{2}\right)^{2}+\Omega^{2}\left(\gamma+\delta\gamma_{q}\right)^{2}}~{},$ (46)
@@
where $\chi$ is referred to as the calibration factor, in units of V / m. This term stems from the fact that most experimental PSDs, calculated from the time-trace recorded on a photodetector, are inherently in units of V2 / Hz. However, given we are concerned with the motion of a nanoparticle, we require a PSD in units of m2 / Hz to determine relevant quantities, such as the position resolution. Formally, $\chi$ is equivalent to $\beta^{\ell,m}_{q}$ from Appendices C and D multiplied by the responsivity (in units of V / W) of the detectors used. For ease of fitting, here we choose to simplify this PSD to one of the form
$S_{qq}(\Omega)$ $=\frac{A}{(B^{2}-\Omega^{2})^{2}+\Omega^{2}C^{2}}~{}.$ (47)
@@
Hence, by fitting Eq. (47) to data taken at thermal equilibrium ($T_{0}=300$ K), we can extract the calibration factor in terms of known quantities. That is,
$\chi=\sqrt{\frac{A}{C}\frac{\pi m}{k_{B}T_{0}}}~{}.$ (48)
Appendix H Temperature estimation